OpenStudy (anonymous):

solve this and show all your work (4,-2) (b,9); d=5squareroot sign5 solve using the distance formula

OpenStudy (anonymous):

Distance \[=5 \sqrt{5}\]

OpenStudy (anonymous):

Distance Formula \[\sqrt{(x_2 -x_1)^2+(y_2 -y_1)^2)}\]

OpenStudy (anonymous):

Plug in what you know and set it equal to \[5\sqrt {5}\]

OpenStudy (anonymous):

i dont understand my book says the answer should be 2,6

OpenStudy (anonymous):

Okay, let work it through... \[\sqrt{(4-b)^2+(-2-9)^2}=5\sqrt{5}\]

OpenStudy (anonymous):

\[\sqrt{(16-8b+b^2)+(-11)^2}=5\sqrt{5}\] (16-8b+b^2)+(-11)^2=125